Alphonse has three times as many marbles as Beatrice. If Alphonse would give 15 of his marbles to Beatrice then he would have twice as many marbles as she does. How many marbles must Alphonse give to Beatrice so that they each have the same number?
(A) 30 (B) 45 (C) 60 (D) 90 (E) Not enough Information

Answer by mananth(16946)

let beatrice have x marbles
Al will have 3x
3x-15 = 2* x
3x-2x=15
x=15
beatrice has 15 marbles
Alphonse has 45 marbles
15+45 / 2 = 30
So alphonse must give 15 marbles to beatrice so that they have the same number 30 marbles